lemon1boy1:

Triangle ABC is similar to triangle DEF. Write the equation, in slope-intercept form, of the side of triangle DEF that is parallel to segment BC. Triangle ABC and DEF are shown. A is at negative 2, 1. B is at negative 2, 4. C is at 1, 1. D is at 4, 2. E is at 4, 4. F is at 6, 2.

surjithayer:

A(-2,1),B((-2,4),C(1,1) D(4,2),E(4,4),F(6,2) slope of BC \[=\frac{ 1-4 }{ 1-(-2) }=\frac{ -3 }{ 3 }=-1\] slope of EF \[=\frac{ 2-4 }{ 6-4 }=\frac{ -2 }{ 2}=-1\] slope of BC=slope of EF=-1 eq. of line EF is \[y-4=-1(x-4),\] \[y-4=-x+4\] or \[y=-x+8\]

lemon1boy1:

I dont need help thank you